OpenStudy (anonymous):

1. (22+5i) + (3+19i) 2. (4+7i) - (5+i)

OpenStudy (anonymous):

Just add/ substract constant with constant and i terms together: (22+5i) + (3+19i) = 25 + 24i (4+7i) - (5+i) = -1+ 6i

OpenStudy (anonymous):

wat about this -> state the values of the a and b components of a + bi form for each complex. state whether the value is the real partor the imaginary part of the number. 1. 3-2i 2. -2-i 3. 3i 4. 0+ 0i 5. -7-4i plzz

OpenStudy (anonymous):

1. a = 2 real , b= -2 imaginary 2. a = -2 real, b = -1 imaginary similary for the rest

OpenStudy (anonymous):

can u show work please how u did it i want to understand

OpenStudy (anonymous):

sorry nvm i understand

OpenStudy (anonymous):

the form is a + bi match with 3 - 2i Can you compare a = ?, b = ?

OpenStudy (anonymous):

wat about 2/3 squarerrot of -162

OpenStudy (anonymous):

(2/3) sqrt ( -162) = (2/3) i *sqrt ( 2 * 9^2) = (2/3) 9i sqrt(2)

OpenStudy (anonymous):

= 6i sqrt (2)

OpenStudy (anonymous):

thankss wat about ( 2+5i)^2